What’s the answers for all 4 questions
Test: Week #1 Exam #2 Sheikcal, Inc provides the fotowing data for the year 2017 Sales Reve Sales Retuns and Allowances Sales Discount 420 000 1400 This Question: 1 pt 417,750 250.000 Cost of Goods Olivera, Inc provides the foilowing data for 2017 he cost of goods sold as a pencentage of net sales ever Round your answer to to dea Sales Revenue Sales Retums and Alowances Sales Discounts Netl Sales Revenue Cost of Goods Sold 627,000 24,000 7.000 505.000 370.000 A.6258% B.6245% C. 6190% The gross pront as a percentage of net salesis(Round your answer to hwo decimal pla A.6200% DB. 3792% C. 39.09% D. 30.04% Horontal anlrovides a year to-year comparison of a companys pertomance in isent pnod iscontinued operations which of the olowing staements is incomedt? O True and losses on the sale of plant assets are no ported as dscontinued oprato discontinued operations heading of the inoe saement ncsts of th es ht have bean sol anallys do not include the esuts of discontnued oporations in francial analysn on discontinued operatiens is ported with an adtion for the picable ncoe PIC.COLLAGE
Answer #1
1) OLIVERA INC:
Gross profit = 596000-370000 = $226000
Gross profit ratio = 226000/596000 = 37.92%
Answer: Option [B]
2) Horizontal analysis…………………………….
*TRUE
3) COGS as a % of net sales = 260000/417750 = 62.24%
Answer: Option [D]
4) Discontinued operations………..
*Gains and losses on the sale of plant assets
are not reported as discontinued operations.
These are reported only when they relate
to a discontinued segment.